What is the propotion of polymorphic loci?What is the propotion of polymorphic loci? Could

Could you explain how the answers in red are the correct answers? I don't understand thr process that was used to get to the correct answer.
Answer #1

Ans 1. a) Polymorphic loci is a locus at which two or more alleles coexist. In the given data table, there are 10 loci shown from A to J. Among these, two loci (C and H) have zero heterozygosity. So, these two loci are not polymorphic. Rest 8 loci are polymorphic to some extent.

So, the proportion of polymorphic loci = 8/10 = 0.8 (correct answer).

Ans 1 b) Average heterozygosity across all loci = sum of heterozygosity at each locus/total number of loci

= (0.48 + 0.03 + 0 + 0.20 + 0.50 + 0.58 + 0.73 + 0 + 0.17 +0.17)/10 = 0.286 (correct answer).

(Note: Please post rest of the questions separately).